# RELATE: A Schema-Agnostic Perceiver Encoder for Multimodal Relational Graphs

### v1

@@ 1 Introduction
To address these challenges, we introduce the RELATE (Relational Encoder for Latent Aggregation of Typed Entities), a schema-agnostic feature encoder designed for heterogeneous graphs with multimodal node features. RELATE uses modality-specific modules shared across all columns of the same type (e.g., categorical, numerical, textual), followed by a Perceiver-style cross-attention layer jaegle2021perceiver that compresses the set of column embeddings into a compact, fixed-size representation per node. This design is permutation-invariant to column order, accommodates varying schemas across node types, and scales to datasets with hundreds of features by attending from a small number of learnable latent queries.
We evaluate RELATE on a wide range of node classification and regression tasks from the RelBench benchmark robinson2024relbench , comparing against schema-specific encoders used in prior RDL frameworks.
@@
Figure 1: Overview of the RELATE architecture. Columns are encoded by shared modality-specific encoders, aggregated with column metadata, and summarized by a cross-attention transformer into fixed-size node embeddings for GNNs.
@@ 3.3 Permutation-Invariant Column Aggregation
To transform a variable-length set of column embeddings into a fixed-size node representation, we adopt a cross-attention module inspired by PerceiverIO jaegle2021perceiver . A shared sequence of $L$ learnable latent tokens $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{L}]\in\mathbb{R}^{L\times d}$ serves as queries, while the column embeddings for node $v$, $\mathbf{X}_{v}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{C_{v}}]\in\mathbb{R}^{C_{v}\times d}$, form the input sequence. Since $L\ll C_{v}$, this formulation reduces computational cost by decoupling self-attention from the number of input columns.
@@ 4 Results
We evaluate RELATE on RelBench robinson2024relbench , a recently introduced benchmark for relational deep learning that spans seven real-world multi-table datasets across domains such as e-commerce, healthcare, and finance. These datasets are structured as heterogeneous temporal graphs with multiple node and edge types and rich, multimodal attributes. Our experiments cover two primary tasks: node classification and node regression. We report Area Under the ROC Curve (AUC) for classification and Mean Absolute Error (MAE) for regression. To ensure a fair comparison, all models are trained using the same splits and optimization protocols defined in the benchmark robinson2024relbench . We integrate RELATE into two widely-used architectures—Heterogeneous Graph Transformer (HGT) hu2020heterogeneous and RelGNN robinson2024relbench —demonstrating compatibility with standard backbones. We compare against the default heteroencoder used in RelBench hu2024pytorch , which uses distinct encoders for each node type and column. RELATE instantiates only a single backbone encoder across all node types and shares weights for each modality encoder across node types.
@@ 4.2 Performance Comparison
RELATE matches or slightly underperforms the RelBench heteroencoder on most tasks, while using a single backbone encoder across all node types and not being explicitly tied to any schema. Additionally, the number of learnable parameters is significantly smaller on several tasks, especially those with a large number of tables and features. For example, RELATE outperforms the comparison method on rel-event user-ignore while using only around 29% of learnable parameters compared to the heteroencoder. We find that RELATE performs generally on par with dataset-specific encoders on regression, however in the rel-f1 task RELATE under performs by a larger margin. We expect this is due to rel-f1 having a small number of train examples and a trade-off of single dataset performance and universality.
@@ 4.3 Parameter Efficiency
Figure 2A compares the number of trainable parameters between schema-specific encoders and RELATE across RelBench datasets (see Table˜3 in Section˜A.3 for exact parameter counts) . RELATE maintains a nearly constant parameter footprint across datasets, since it uses a fixed set of modality-specific encoders shared across node types. In contrast, the parameter count of schema-specific encoders grows rapidly with the number of tables and features. As a result, RELATE achieves up to a 5× reduction in parameters on feature-rich datasets such as rel-trial and rel-event, while remaining competitive in cases with fewer features. This stability makes RELATE particularly advantageous for real-world schemas involving hundreds or thousands of attributes, where schema-specific encoders become prohibitively large.
Figure 2: (A) Parameter comparison across RelBench datasets. We compare the number of trainable parameters between schema-specific encoders (Std) and RELATE. Schema-specific encoders grow with the number of tables and features, whereas RELATE maintains a nearly constant footprint. Percentages above the RELATE bars indicate the parameter ratio relative to schema-specific encoders. (B) Attention ablation for RelGNN, reporting average AUC performance relative to cross-attention. Bars show the performance of self-attention compared to cross-attention within each dataset.
@@ 4.4 Ablation Study
We evaluate the role of cross-attention by replacing it with a full self-attention (SA) mechanism applied over the input tokens. As shown in Figure 2B, we report the relative performance of self-attention for RelGNN, averaged over all classification tasks. This substitution yields marginal gains on a few datasets but does not consistently improve performance, with self-attention achieving only 90–101% of cross-attention on average. At the same time, the computational overhead of full self-attention is substantially higher due to its quadratic complexity. In contrast, the cross-attention bottleneck in RELATE provides a more efficient and scalable alternative, preserving accuracy while enabling training on large and feature-rich graphs. Comprehensive AUC and MAE results for both RelGNN and HGT are reported in Appendix A.2.
@@ 6 Conclusion
7 Acknowledgments
We thank Alexandre Dorais, Jialin Dong, Viswa Ganapathy, Dinesh Katupputhur, Prasanna Lalingkar, Karan Paresh, Andrew Pouret, Afreen Shaikh, and Jay Shinigari for their valuable assistance in conducting the experiments.
@@ A.1.3 RELATE Cross-Attention
A.2 Ablation Results
@@
A.3 Parameter Comparison
Table 3 reports the exact parameter counts corresponding to Figure 2. As shown, the parameter footprint of schema-specific encoders increases substantially with the number of tables and features in each dataset, ranging from $1.1\text{M}$ on rel-amazon to over $7.3\text{M}$ on rel-trial. In contrast, RELATE remains nearly constant at $\sim$1.4M parameters across all datasets, since it reuses a fixed set of modality-specific encoders. This property results in significant parameter savings—up to a 5× reduction on feature-rich datasets such as rel-trial and rel-event—while still maintaining competitive performance.

### v2

@@ 1 Introduction
To address these challenges, we introduce the RELATE (Relational Encoder for Latent Aggregation of Typed Entities), a schema-agnostic feature encoder designed for heterogeneous graphs with multimodal node features. RELATE uses modality-specific modules shared across all columns of the same type (e.g., categorical, numerical, textual), followed by a Perceiver-style cross-attention layer jaegle2021perceiver ; graphfm that compresses the set of column embeddings into a compact, fixed-size representation per node. This design is permutation-invariant to column order, accommodates varying schemas across node types, and scales to datasets with hundreds of features by attending from a small number of learnable latent queries.
We evaluate RELATE on a wide range of node classification and a subset of regression tasks from the RelBench benchmark robinson2024relbench , comparing against schema-specific encoders used in prior RDL frameworks.
@@
Figure 1: Overview of the RELATE architecture. Columns are encoded by shared modality-specific encoders, aggregated with column metadata, and summarized by a cross-attention transformer into fixed-size node embeddings for GNNs. Please refer to Appendix A.1.2 for details on how column metadata is aggregated with cell embeddings.
@@ 3.3 Permutation-Invariant Column Aggregation
To transform a variable-length set of column embeddings into a fixed-size node representation, we adopt a cross-attention module inspired by PerceiverIO jaegle2021perceiver ; graphfm . A shared sequence of $L$ learnable latent tokens $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{L}]\in\mathbb{R}^{L\times d}$ serves as queries, while the column embeddings for node $v$, $\mathbf{X}_{v}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{C_{v}}]\in\mathbb{R}^{C_{v}\times d}$, form the input sequence. Since $L\ll C_{v}$, this formulation reduces computational cost by decoupling self-attention from the number of input columns.
@@ 4 Results
We evaluate RELATE on RelBench robinson2024relbench , a recently introduced benchmark for relational deep learning that spans seven real-world multi-table datasets across domains such as e-commerce, healthcare, and finance. These datasets are structured as heterogeneous temporal graphs with multiple node and edge types and rich, multimodal attributes. Our experiments cover two primary tasks: node classification and node regression. We report Area Under the ROC Curve (AUC) for classification and Mean Absolute Error (MAE) for regression. To ensure a fair comparison, all models are trained using the same splits and optimization protocols defined in the benchmark robinson2024relbench . We integrate RELATE into two widely-used architectures—Heterogeneous Graph Transformer (HGT) hu2020heterogeneous and RelGNN robinson2024relbench —demonstrating compatibility with standard backbones (refer to Appendix A.2 for more details). We compare against the default heteroencoder used in RelBench hu2024pytorch , which uses distinct encoders for each node type and column. RELATE instantiates only a single backbone encoder across all node types and shares weights for each modality encoder across node types.
@@ 4.2 Performance Comparison
RELATE matches or slightly underperforms the RelBench heteroencoder on most tasks, while using a single backbone encoder across all node types and not being explicitly tied to any schema. Additionally, the number of learnable parameters is significantly smaller on several tasks, especially those with a large number of tables and features (Figure 2B). For example, RELATE outperforms the comparison method on rel-event user-ignore while using only around 29% of learnable parameters compared to the heteroencoder. We find that RELATE performs generally on par with dataset-specific encoders on regression, however in the rel-f1 task RELATE under performs by a larger margin. We expect this is due to rel-f1 having a small number of train examples and a trade-off of single dataset performance and universality.
Figure 2: (A) Parameter comparison across RelBench datasets. We compare the number of trainable parameters between schema-specific encoders (Std) and RELATE. Schema-specific encoders grow with the number of tables and features, whereas RELATE maintains a nearly constant footprint. Percentages above the RELATE bars indicate the parameter ratio relative to schema-specific encoders. (B) Performance v/s parameter count (log scale) for the HGT backbone. Each line connects a schema-specific encoder (blue) to its RELATE counterpart (green), illustrating the change in both performance and model size. RELATE achieves comparable AUC while using up to 5× fewer parameters.
@@ 4.3 Parameter Efficiency
Figure 2A compares the number of trainable parameters between schema-specific encoders and RELATE across RelBench datasets (see Table˜3 in Section˜A.4 for exact parameter counts) . RELATE maintains a nearly constant parameter footprint across datasets, since it uses a fixed set of modality-specific encoders shared across node types. In contrast, the parameter count of schema-specific encoders grows rapidly with the number of tables and features. As a result, RELATE achieves up to a 5× reduction in parameters on feature-rich datasets such as rel-trial and rel-event, while remaining competitive in cases with fewer features. This stability makes RELATE particularly advantageous for real-world schemas involving hundreds or thousands of attributes, where schema-specific encoders become prohibitively large. Figure 2B further highlights the trade-off between performance and parameter efficiency. In datasets such as rel-hm and rel-avito, RELATE maintains similar performance while using significantly fewer parameters.
Figure 3: Ablation: Attention ablation for RelGNN, reporting average AUC performance relative to cross-attention. Bars show the performance of self-attention compared to cross-attention within each dataset.
@@ 4.4 Ablation Study
We evaluate the role of cross-attention by replacing it with a full self-attention (SA) mechanism applied over the input tokens. As shown in Figure 3, we report the relative performance of self-attention for RelGNN, averaged over all classification tasks. This substitution yields marginal gains on a few datasets but does not consistently improve performance, with self-attention achieving only 90–101% of cross-attention on average. At the same time, the computational overhead of full self-attention is substantially higher due to its quadratic complexity. In contrast, the cross-attention bottleneck in RELATE provides a more efficient and scalable alternative, preserving accuracy while enabling training on large and feature-rich graphs. Comprehensive AUC and MAE results for both RelGNN and HGT are reported in Appendix A.3.
@@ 6 Conclusion
7 Acknowledgments and Disclosure of Funding
We thank Alexandre Dorais, Jialin Dong, Viswa Ganapathy, Dinesh Katupputhur, Prasanna Lalingkar, Karan Paresh, Andrew Pouret, Afreen Shaikh, and Jay Shinigari for their valuable assistance in conducting the experiments and Qiunan Ma for helping with the figures. This project was also supported by NSF CAREER Award RI:2146072, NSF award CIF:RI:2212182 as well as generous gifts from the CIFAR Azrieli Global Scholars Program (D.L and E.L.D).
@@ A.1.3 RELATE Cross-Attention
A.2 Backbone Models
RelGNN:
The Relational Graph Neural Network (RelGNN) robinson2024relbench serves as a strong heterogeneous GNN baseline that extends message passing to relational databases represented as multi-entity graphs. Unlike traditional GNNs that operate on homogeneous graph structures, RelGNN performs message passing over relational graphs where nodes represent entities (e.g., tables or records) and edges encode typed relationships derived from foreign key links. This schema-aware formulation enables the model to jointly reason over multiple interconnected tables while preserving the structural and semantic integrity of the underlying database. In our experiments, we adopt the default hyperparameter configuration provided in RelBench robinson2024relbench .
HGT:
The Heterogeneous Graph Transformer (HGT) hu2020heterogeneous extends the transformer architecture to handle multi-relational graphs with multiple node and edge types. Unlike standard graph transformers that assume homogeneous structures, HGT introduces type-specific transformation matrices and relation-aware attention to model heterogeneity in both message passing and aggregation. Each edge type is associated with a unique attention function, allowing the model to capture asymmetric relationships between different entity types.
In our experiments, we use the HGT+PE variant, which augments the standard HGT with Laplacian positional encodings (LapPE) arora2025exploiting computed over sampled subgraphs to incorporate structural context. We employ the HGTConv implementation from PyTorch Geometric fey2024graph, integrated into the RDL pipeline robinson2024relbench . The model consists of 2 HGT layers with 4 attention heads each, residual connections, and layer normalization. The dimensionality of the LapPE is set to 4 for most datasets and 2 for datasets such as rel-amazon and rel-hm.
A.3 Ablation Results
@@
A.4 Parameter Comparison
Table 3 reports the exact parameter counts corresponding to Figure LABEL:fig:ablation. As shown, the parameter footprint of schema-specific encoders increases substantially with the number of tables and features in each dataset, ranging from $1.1\text{M}$ on rel-amazon to over $7.3\text{M}$ on rel-trial. In contrast, RELATE remains nearly constant at $\sim$1.4M parameters across all datasets, since it reuses a fixed set of modality-specific encoders. This property results in significant parameter savings—up to a 5× reduction on feature-rich datasets such as rel-trial and rel-event—while still maintaining competitive performance.